Let Z be a standard normal random variable.

P(c<=Z<=-0.66)=0.2388

Carry your intermediate computations to at least four decimal places. Round your answer to two decimal places.
Answer #1

2 Let be standard normal randons variable ZN(0,0) PC cez _ -0.66) = 0.2388 P(Z <-6-66) - P(Z¢) = 6-2388 0.2546 - P(Z <C) = 6